OpenStudy (anonymous):

To find an object's density (D), first measure its mass (m) and volume (V). Then use the following equation: D = m/V Densities of Common Substances Substance Density (g/cm3) Substance Density (g/cm3) Helium (gas) 0.0001663 Copper (solid) 8.96 Iron pyrite (solid) 5.02 Silver (solid) 10.50 Water (liquid) 1.00 Lead (solid) 11.35 Examine the above equation and table. Robin has a lead ball with a mass of 454 g. What is its volume? 476.7 g3 40 g3 40 cm3 11.35 g/cm3

OpenStudy (anonymous):

It has to be B or C so

OpenStudy (agent0smith):

\[\Large D= \frac{ m }{ V }\] you have the mass, 454. The density of lead is whatever it is in the table. Plug in numbers and rearrange to find V. What is volume measured in? cubic grams, or cubic centimetres?

OpenStudy (anonymous):

cubic grams

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I would say 40 g3 because I did 11.35 / 454 and the answer was 40.

OpenStudy (anonymous):

Sir?

OpenStudy (agent0smith):

Grams is a measurement for weight, and cubic grams isn't anything. Volume is measured in cubic metres, or centimetres etc. \[\Large 11.35= \frac{ 454 }{ V }\] so \[\Large V = \frac{ 454 }{ 11.35 }\]

OpenStudy (anonymous):

The answer will be 40cm3

OpenStudy (anonymous):

ohh nevermind 11.35 g/cm3 is the answer

OpenStudy (agent0smith):

11.35 g/cm3 that is the density of lead...

OpenStudy (agent0smith):

Volume is measured in cm^3.

OpenStudy (anonymous):

So 11.35 / 454 = 40 cm^3

OpenStudy (agent0smith):

11.35/454 = 0.025 and is not correct. You want : \[\Large V = \frac{ 454 }{ 11.35 } \]

OpenStudy (anonymous):

I meant 454 / 11.35= 40cm^3

OpenStudy (anonymous):

That's what I am saying.

OpenStudy (agent0smith):

If you'd written it once i might've understood... but twice you said that 11.35 / 454 = 40 so i couldn't tell if you actually knew what you were calculating.

OpenStudy (anonymous):

Oh I am sorry. I will correct that in the future thank you